OpenStudy (anonymous):

Slope 3, (2, 6) how to change to slope intercept form

OpenStudy (solomonzelman):

y=mx+b m=slope b=y-intercept So you know that y=3x+b but the y-intercept is unknown, plug in your point (6)=3(2)+b 6=6+b b=0 So it's going to be just y=3x